# Lighter, Better, Faster Multi-Source Domain Adaptation with Gaussian Mixture Models and Optimal Transport

### v2

@@ Abstract
In this paper, we tackle Multi-Source Domain Adaptation (MSDA), a task in transfer learning where one adapts multiple heterogeneous, labeled source probability measures towards a different, unlabeled target measure. We propose a novel framework for MSDA, based on Optimal Transport (OT) and Gaussian Mixture Models (GMMs). Our framework has two key advantages. First, OT between GMMs can be solved efficiently via linear programming. Second, it provides a convenient model for supervised learning, especially classification, as components in the GMM can be associated with existing classes. Based on the GMM-OT problem, we propose a novel technique for calculating barycenters of GMMs. Based on this novel algorithm, we propose two new strategies for MSDA: GMM-WBT and GMM-DaDiL. We empirically evaluate our proposed methods on four benchmarks in image classification and fault diagnosis, showing that we improve over the prior art while being faster and involving fewer parameters.
@@ 1 Introduction
In the context of DA, a prominent framework is Optimal Transport (OT) [5], which is a field of mathematics concerned with the displacement of mass at least effort. This theory has been applied for DA in multiple ways, especially by (i) mapping samples between domains [6] and (ii) learning invariant representations [7]. For MSDA, OT has been used for aggregating the multiple source domains into a barycentric domain [8, 9], which is later transported to the target domain, or by weighting source domain measures [10].
This work considers the problem of Wasserstein Dictionary Learning (WDL), which was initially proposed by [11] for histogram data. This problem was later generalized by [12], for empirical measures, which allowed its application to MSDA. In [12], one learns to express domains in MSDA as a barycenter of atom measures, which have a free, learnable support. As a result, the work of [12] learns how to interpolate distributional shift between the measures in MSDA.
However, previous algorithms relying on Wasserstein barycenters, such as Wasserstein Barycenter Transport (WBT) [8, 9] and Dataset Dictionary Learning (DaDiL) [12], are limited in scale, since the number of points the support of the empirical measures scale with the number of samples in the original datasets. As a consequence, previous works such as [8, 9] are limited to small scale datasets, or rely on mini-batch optimization [12], which introduces artifacts in the OT. To tackle the aforementioned limitations, in this paper we propose a novel, parametric framework for barycentric-based MSDA algorithms. based on recent developments in OT between Gaussian Mixture Models (GMMs) [13]. We present an overview of our methods in figure 1.
@@
Figure 1: Overview of proposed methods. Squares represent actual domains, circles represent barycenters and triangles represent learned measures. Blue and orange elements represent labeled and unlabeled measures respectively. In GMM-WBT, a labeled GMM is determined for the target domain by transporting the barycenter of sources. In GMM-DaDiL, we learn to express each domain as a barycenter of learned GMMs, called atoms, through dictionary learning.
Summary of Contributions. Our contributions are threefold: 1. We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 1); 2. We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3. We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4). We highlight that, while GMMs were previously employed in single source DA [14, 15], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
Paper Organization. The rest of this paper is divided as follows. Section 2 covers the background behind our method. Section 3 covers our methodological contributions. Section 4 explores the empirical validation of our method with respect other OT-based MSDA algorithms, where we show that our methods significantly outperform prior art. Finally, section 5 concludes this paper.
@@ 2 Preliminaries
In this section, we cover the necessary background for our paper. We divide our discussion into Gaussian mixtures, domain adaptation, optimal transport and dataset dictionary learning.
Gaussian Mixtures. We denote the set of probability measures over a set $\mathcal{X}$ as $\mathbb{P}(\mathcal{X})$. A Gaussian measure corresponds to $P_{\theta}\in\mathbb{P}(\mathcal{X})$ with the following density,
@@
where $\Delta_{K}=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k=1}^{K}p_{k}=1\}$. Following [13], we denote the subset of $\mathbb{P}(\mathbb{R}^{d})$ of probability measures which can be written as Gaussian mixtures with less than $K$ components by $\text{GMM}_{d}(K)$, and $\text{GMM}_{d}(\infty)=\cup_{k\geq 0}\text{GMM}_{d}(K)$.
@@
where $\Theta=\{\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)}\}_{k=1}^{K}:\sum_{k}p_{k}=1\text{ and }\mathbf{C}_{k}^{(P)}\in\mathbb{S}^{d}\}$. While equation 2 has no closed-form solution, one can solve this optimization problem through the celebrated Expectation-Maximization (EM) algorithm [16].
Learning Theory and Domain Adaptation. In this paper, we focus on the problem of classification. Given a feature space $\mathcal{X}=\mathbb{R}^{d}$ and a label space $\mathcal{Y}=\{1,\cdots,n_{cl}\}$, this problem corresponds to finding $h\in\mathcal{H}\subset\mathcal{X}^{\mathcal{Y}}$ that correctly classifies data $\{(\mathbf{x}_{i}^{(Q)},y_{i}^{(Q)})\}_{i=1}^{n}$.
In this paper, we use the Empirical Risk Minimization (ERM) framework [17], as it is useful for domain adaptation theory. As follows, one assumes $\mathbf{x}_{i}^{(Q)}\overset{iid}{\sim}Q$, for a measure $Q\in\mathbb{P}(\mathcal{X})$, and $h_{0}\in\mathcal{X}^{\mathcal{Y}}$ such that $y_{i}^{(Q)}=h_{0}(\mathbf{x}_{i}^{(Q)})$. $h_{0}$ is called ground-truth labeling function. Given a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, a classifier may be defined through risk minimization, i.e., $h^{\star}=\text{argmin}_{h\in\mathcal{H}}\mathcal{R}_{Q}(h)$, for $\mathcal{R}_{Q}(h)=\mathbb{E}_{Q}[\mathcal{L}(h(\mathbf{x}),h_{0}(\mathbf{x}))]$. This strategy is oftentimes impractical as $Q$ and $h_{0}$ are unknown. As a result, one resorts to the minimization of the empirical risk, i.e., $\hat{h}=\text{argmin}_{h\in\mathcal{H}}\hat{\mathcal{R}}_{Q}(h)$, where ^$\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(h(\mathbf{x}_{i}^{(Q)}),y_{i}^{(Q)})$.
From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [17, Section 6]. As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
Nevertheless, the assumption that unseen examples come from a fixed measure $P$ is seldom verified in practice [1], since the conditions upon which data is acquired may change. In this case, models are required to adapt to new data, but at the same time re-training a model from the scratch is likely costly and data intensive. A solution consists of using transfer learning [3], in which one re-uses knowledge from a source domain or task to facilitate the learning on a target domain or task.
@@
where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$. Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$. A more tractable formulation was proposed by Kantorovich [18, Section 2.3.], and relies on OT plans,
@@
$\mathcal{W}_{2}^{2}(P,Q)=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\text{Tr}\biggr{(}\mathbf{C}^{(P)}+\mathbf{C}^{(Q)}-2(\mathbf{S}^{(P)}\mathbf{C}^{(Q)}\mathbf{S}^{(P)})^{\nicefrac{{1}}{{2}}}\biggr{)},$
@@
$\mathcal{W}_{2}^{2}(P,Q)$ $=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\lVert\mathbf{s}^{(P)}-\mathbf{s}^{(Q)}\rVert_{2}^{2}.$ (6)
@@
In this paper, we use the GMM-OT framework of [13], which is convenient to our setting for 2 reasons. First, they are able to represent measures with sub-populations, such as those commonly encountered in classification and domain adaptation. Second, they yield a tractable OT problem, when $\gamma$ is further restricted to be a GMM itself, that is,
@@
Dataset Dictionary Learning [12] is an OT-based framework for expressing probability measures as barycenters of synthetic measures, called atoms. The framework is inspired by dictionary learning literature [11]. The authors introduce atoms $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ and barycentric coordinates $\Lambda=\{\lambda_{\ell}\}_{\ell=1}^{N}$, such that each measure in MSDA is expressed as a Wasserstein barycenter $\hat{Q}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$. This framework leads to the following optimization problem,
$\mathcal{L}(\Lambda,\mathcal{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{W}_{2}(\hat{Q}_{T},\mathcal{B}(\lambda_{\ell};\mathcal{P}))+\sum_{\ell=1}^{N}\mathcal{W}_{c,2}(\hat{Q}_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2},$ (9)
where $c\biggr{(}(\mathbf{x}^{(Q_{\ell})},\mathbf{y}^{(Q_{\ell})}),(\mathbf{x}^{(B_{\ell})},\mathbf{y}^{(B_{\ell})})\biggr{)}=\lVert\mathbf{x}^{(Q_{\ell})}-\mathbf{x}^{(B_{\ell})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}^{(Q_{\ell})}-\mathbf{y}^{(B_{\ell})}\rVert_{2}^{2}$, and $\beta\geq 0$ is a constant expressing how costly it is to move samples from different classes. This framework makes it easier to express the distributional shift between the different $\mathcal{Q}=\{\hat{Q}_{1},\cdots,\hat{Q}_{N_{S}},\hat{Q}_{T}\}$. Especially, since $\hat{B}_{\ell}=\mathcal{B}(\lambda_{\ell};\mathcal{P})$ is labeled, one can synthesize labeled target domain data by reconstructing the target measure with $\lambda_{T}:=\lambda_{N+1}$.
@@ 3.1 First Order Analysis of $\mathcal{MW}_{2}$
In this section we analyze $P\mapsto\mathcal{MW}_{2}(P,Q)$, for a fixed $Q$. We are particularly interested on how to map the components of $P$ towards $Q$, while minimizing this distance. The following theorem provides us a strategy, [^1] We refer readers to our supplementary materials for all proofs and derivations.
@@
$\mathbf{m}_{i}=T_{\omega^{\star}}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega^{\star}}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)},$ (10)
@@ 3.2 Supervised Mixture-Wasserstein Distances
In this paper, we consider supervised learning problems. As such, it is necessary to equip the components of GMMs with labels that represent the classes in the datasets. We propose doing so through a simple heuristic, based on the law of total probability,
$P(y|k)$ $=\sum_{i=1}^{n}P_{\theta}(\mathbf{x}_{i}|k)P(y|\mathbf{x}_{i})=\sum_{i=1}^{n}\biggr{(}\dfrac{p_{k}P_{k}(\mathbf{x}_{i})}{\sum_{k^{\prime}}p_{k^{\prime}}P_{k^{\prime}}(\mathbf{x}_{i})}\biggr{)}P(y|\mathbf{x}_{i}),$ (11)
where $P_{\theta}(\mathbf{x}_{i}|k)$ corresponds to the probability that $\mathbf{x}_{i}$ comes form the $k-$th component of the GMM. Henceforth, we denote the soft-labels of components by $\mathbf{y}_{i}^{(P)}$. We may use the labeled GMM for obtaining a classifier, by sampling $\mathbf{x}\sim P_{\theta}$. If $\mathbf{x}$ is sampled from the $k-$th component, we simply assign $P(y|k)$ as its soft-label. We illustrate our strategy in figure 2.
(a) Gaussian Mixture $P_{\theta}$
(b) $\text{argmax}_{k}P_{\theta}(\mathbf{x}|k)$.
(c) Labeled GMM
Figure 2: Illustration of the labeling process of Gaussian mixtures. (a) First, the mixture $P_{\theta}$ is learned. (b) Then, one estimates $P_{\theta}(\mathbf{x}|k)$ for each $\mathbf{x}_{i}$. (c) The label of each component can be estimated using the law of total probability.
@@
$\mathcal{SMW}_{2}(P,Q)^{2}$ $=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{argmin}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}(\mathcal{W}_{2}(P_{i},Q_{j})^{2}+\beta\lVert\mathbf{y}_{i}^{(P)}-\mathbf{y}_{j}^{(Q)}\rVert_{2}^{2}).$ (12)
@@
Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 10. Furthermore, for $\mathbf{y}_{i}$,
$\hat{\mathbf{y}}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (13)

### v3

@@ Abstract
[…]
@@ 1 Introduction
In the context of DA, a prominent framework is Optimal Transport (OT) [5], which is a field of mathematics concerned with the displacement of mass at least effort. This theory has been applied for DA in multiple ways, especially by (i) mapping samples between domains [6] and (ii) learning invariant representations [7]. For MSDA, OT has been used for aggregating the multiple source domains into a barycentric domain [8, 9], which is later transported to the target domain, or by weighting source domain measures [10]. Our work considers the problem of Wasserstein Dictionary Learning (WDL), initially proposed by [11] for histogram data. This problem was later generalized by [12], for empirical measures, which allowed its application to MSDA. In [12], one expresses domains in MSDA as a barycenter of atom measures, which have a free, learnable support. As a result, the work of [12] learns how to interpolate distributional shift between the measures in MSDA.
[…]
@@
[…]
However, previous algorithms relying on Wasserstein barycenters, such as WBT [8, 9] and DaDiL [12], are limited in scale, since the number of points the support of the empirical measures scale with the number of samples in the original datasets. As a consequence, previous works such as [8, 9] are limited to small scale datasets, or rely on mini-batch optimization [12], which introduces artifacts in the OT. To tackle these limitations, in this paper we propose a novel, parametric framework for barycentric-based MSDA algorithms. based on OT between GMMs [13]. We present an overview of our methods in figure 1.
Our contributions are threefold: 1. We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 1); 2. We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3. We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4). We highlight that, while GMMs were previously employed in single source DA [14, 15], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
The rest of this paper is divided as follows. Section 2 covers the background behind our method. Section 3 covers our methodological contributions. Section 4 explores the empirical validation of our method with respect other OT-based MSDA algorithms, where we show that our methods significantly outperform prior art. Finally, section 5 concludes this paper.
@@ 2 Preliminaries
2.1 Gaussian Mixtures
We denote the set of probability measures over a set $\mathcal{X}$ as $\mathbb{P}(\mathcal{X})$. A Gaussian measure corresponds to $P_{\theta}\in\mathbb{P}(\mathcal{X})$ with density,
@@ 2.1 Gaussian Mixtures
where $\Delta_{K}=\{\mathbf{p}\in\mathbb{R}^{K}_{+}:\sum_{k=1}^{K}p_{k}=1\}$. Following [13], we denote the subset of $\mathbb{P}(\mathbb{R}^{d})$ of probability measures which can be written as Gaussian mixtures with less than $K$ components by $\text{GMM}_{d}(K)$, and $\text{GMM}_{d}(\infty)=\cup_{k\geq 0}\text{GMM}_{d}(K)$.
@@
where $\Theta=\{\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{C}_{k}^{(P)}\}_{k=1}^{K}:\mathbf{m}_{k}^{(P)}\in\mathbb{R}^{d}\text{ and }\mathbf{C}_{k}^{(P)}\in\mathbb{S}^{d}\}$. While equation 2 has no closed-form solution, one can solve this optimization problem through the celebrated Expectation-Maximization (EM) algorithm [16].
2.2 Domain Adaptation
In this paper, we focus on the problem of classification. Given a feature space $\mathcal{X}=\mathbb{R}^{d}$ and a label space $\mathcal{Y}=\{1,\cdots,n_{cl}\}$, this problem corresponds to finding $h\in\mathcal{H}\subset\mathcal{Y}^{\mathcal{X}}$ that correctly classifies data $\{(\mathbf{x}_{i}^{(Q)},y_{i}^{(Q)})\}_{i=1}^{n}$.
We use the Empirical Risk Minimization (ERM) framework [17], as it is useful for domain adaptation theory. As follows, one assumes $\mathbf{x}_{i}^{(Q)}\overset{iid}{\sim}Q$, for a measure $Q\in\mathbb{P}(\mathcal{X})$, and $h_{0}:\mathcal{X}\rightarrow\mathcal{Y}$ such that $y_{i}^{(Q)}=h_{0}(\mathbf{x}_{i}^{(Q)})$. $h_{0}$ is called ground-truth labeling function. Given a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, a classifier may be defined through risk minimization, i.e., $h^{\star}=\text{argmin}_{h\in\mathcal{H}}\mathcal{R}_{Q}(h)$, for $\mathcal{R}_{Q}(h)=\mathbb{E}_{Q}[\mathcal{L}(h(\mathbf{x}),h_{0}(\mathbf{x}))]$. This strategy is oftentimes impractical as $Q$ and $h_{0}$ are unknown. As a result, one resorts to the minimization of the empirical risk, i.e., $\hat{h}=\text{argmin}_{h\in\mathcal{H}}\hat{\mathcal{R}}_{Q}(h)$, where ^$\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(h(\mathbf{x}_{i}^{(Q)}),y_{i}^{(Q)})$.
From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [17, Section 6]. As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $Q$. Nevertheless, the assumption that unseen examples come from a fixed measure $Q$ is seldom verified in practice [1], since the conditions upon which data is acquired may change. In this case, models are required to adapt to new data, but at the same time re-training a model from the scratch is likely costly and data intensive. A solution consists of using transfer learning [3], in which one re-uses knowledge from a source domain or task to facilitate the learning on a target domain or task.
@@ 2.3 Optimal Transport
where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$, and $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a ground-cost, that is, a measure of transportation effort. Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$. A more tractable formulation was proposed by Kantorovich [18, Section 2.3.], and relies on OT plans,
@@
$\mathcal{W}_{2}(P,Q)^{2}=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\text{Tr}\biggr{(}\mathbf{C}^{(P)}+\mathbf{C}^{(Q)}-2(\mathbf{S}^{(P)}\mathbf{C}^{(Q)}\mathbf{S}^{(P)})^{\nicefrac{{1}}{{2}}}\biggr{)},$
@@
$\mathcal{W}_{2}(P,Q)^{2}$ $=\lVert\mathbf{m}^{(P)}-\mathbf{m}^{(Q)}\rVert^{2}_{2}+\lVert\mathbf{s}^{(P)}-\mathbf{s}^{(Q)}\rVert_{2}^{2}.$ (6)
@@
Remark. Here, we give further insight into the hypothesis of using axis-aligned Gaussian measures. We use this assumption for numerical stability purposes, i.e., estimating the covariance matrix of GMMs in high dimensions is much more difficult than estimating the standard deviation vector $\mathbf{s}^{(P)}$. Here, one has two choices. First, it is possible to introduce a transformation so as to force features to be uncorrelated (e.g., through principal components analysis). This approach, nonetheless, requires more data points per domain than features, which is not always feasible. Conversely, one can increase the number of components for expressing the shape of the data (see Fig. 2). As we show in our experiments sections, we achieve good adaptation performance, even while sampling points from axis-aligned GMMs.
Figure 2: Illustration of axis-aligned GMMs. This hypothesis leads to GMMs that need more components to express the underlying data distribution.
We use the GMM-OT framework of [13], which is convenient to our setting for 2 reasons. First, they are able to represent measures with sub-populations, such as those commonly encountered in classification and domain adaptation. Second, they yield a tractable OT problem, when $\gamma$ is further restricted to be a GMM itself, that is,
@@
DaDiL [12] is an OT-based framework for expressing probability measures as barycenters of synthetic measures, called atoms. In this case, the authors use empirical measures, i.e., $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{(\mathbf{x}_{i}^{(P)},\mathbf{y}_{i}^{(P)})}$. The framework is inspired by dictionary learning literature [11]. The authors introduce atoms $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ and barycentric coordinates $\Lambda=\{\lambda_{\ell}\}_{\ell=1}^{N}$, such that each measure in MSDA is expressed as a Wasserstein barycenter $\hat{Q}_{\ell}=\mathcal{B}(\lambda_{\ell},\mathcal{P})$. This framework leads to the following optimization problem,
$(\Lambda^{\star},\mathcal{P}^{\star})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{W}_{2}(\hat{Q}_{T},\mathcal{B}(\lambda_{\ell};\mathcal{P}))+\sum_{\ell=1}^{N}\mathcal{W}_{c,2}(\hat{Q}_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2},$ (9)
where $c\biggr{(}(\mathbf{x}^{(Q_{\ell})},\mathbf{y}^{(Q_{\ell})}),(\mathbf{x}^{(B_{\ell})},\mathbf{y}^{(B_{\ell})})\biggr{)}=\lVert\mathbf{x}^{(Q_{\ell})}-\mathbf{x}^{(B_{\ell})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}^{(Q_{\ell})}-\mathbf{y}^{(B_{\ell})}\rVert_{2}^{2}$, and $\beta\geq 0$ is a constant expressing how costly it is to move samples from different classes. This framework makes it easier to express the distributional shift between the different $\mathcal{Q}=\{\hat{Q}_{1},\cdots,\hat{Q}_{N_{S}},\hat{Q}_{T}\}$. Especially, since $\hat{B}_{\ell}=\mathcal{B}(\lambda_{\ell},\mathcal{P})$ is labeled, one can synthesize labeled target domain data by reconstructing the target measure with $\lambda_{T}:=\lambda_{N+1}$.
@@ 3.1 First Order Analysis of $\mathcal{MW}_{2}$
In this section we analyze $P\mapsto\mathcal{MW}_{2}(P,Q)^{2}$, for a fixed $Q$. We are particularly interested on how to map the components of $P$ towards $Q$, while minimizing this distance. The following theorem provides us a strategy,
@@
$\hat{\mathbf{m}}_{i}=T_{\omega^{\star}}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\hat{\mathbf{s}}_{i}=T_{\omega^{\star}}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)},$ (10)
@@
Proof.
Our proof relies on the analysis of $\{(\mathbf{m}_{i},\mathbf{s}_{i})\}_{i=1}^{K_{P}}\mapsto\mathcal{MW}_{2}(P,Q)^{2}$ (c.f., equation 8). Given $\omega^{\star}=\text{GMMOT}(P,Q)$,
and, by equating this last term to 00, one gets the desired equality. ∎
@@ 3.2 Supervised Mixture-Wasserstein Distances
In this paper, we consider supervised learning problems. As such, it is necessary to equip the components of GMMs with labels that represent the classes in the datasets. We propose doing so through a simple heuristic, especially, we model $P(\mathbf{x}|y)$ through a GMM. We then concatenate the $n_{c}$ obtained GMMs, and assign, for the $k-$th GMM of the $y-$th class, $v_{k,y^{\prime}}^{(P)}=\delta(y^{\prime}-y)$, i.e., a vector of $n_{c}$ components, and $1$ on the $y-$th entry. We can assure that the resulting weights sum to $1$ by dividing their value by $\sum_{y=1}^{n_{c}}\sum_{k=1}^{K}p_{k,y}$, where $p_{k,y}$ corresponds to the weight of the $k-$th component of the $y-$th GMM.
Given a GMM $\{p_{k},\mathbf{m}_{k}^{(P)},\mathbf{s}_{k}^{(P)},\mathbf{v}_{k}^{(P)}\}_{k=1}^{K}$, we define a classifier through Maximum a Posteriori (MAP) estimation. This strategy is carried out through,
$\hat{h}_{MAP}(\mathbf{x})=\underset{y=1,\cdots,n_{c}}{\text{argmax}}\,P(y|\mathbf{x})=\sum_{k=1}^{K}\underbrace{P_{\theta}(k|\mathbf{x})}_{p_{k}P_{k}(\mathbf{x})/\sum_{k^{\prime}}p_{k^{\prime}}P_{k^{\prime}}(\mathbf{x})}\underbrace{P(y|k)}_{v_{k,y}^{(P)}},$ (11)
we use this classifier in a few illustrative examples in section 4.3.
Remark. In equation 11, we are implicitly assuming that the component $k$ is conditionally independent with $y$ given $\mathbf{x}$. This remark is intuitive, as $\mathbf{x}$ explains, at the same time, the component and the label.
@@
$\mathcal{SMW}_{2}(P,Q)^{2}$ $=\underset{\omega\in\Gamma(\mathbf{p},\mathbf{q})}{\text{min}}\,\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}(\mathcal{W}_{2}(P_{i},Q_{j})^{2}+\beta\lVert\mathbf{v}_{i}^{(P)}-\mathbf{v}_{j}^{(Q)}\rVert_{2}^{2}).$ (12)
@@
Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{v}_{i}^{(P)}$ and $\mathbf{v}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 10. Furthermore, for $\mathbf{v}_{i}$,
$\hat{\mathbf{v}}_{i}=T_{\omega}(\mathbf{v}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{v}_{j}^{(Q)}.$ (13)
Proof.
The label distance term in $\mathcal{SMW}$ is independent of $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$, hence the optimality conditions of these variables remain unchanged. Therefore, the first-order optimality condition with respect $\mathbf{v}_{i}$ is,
$\dfrac{\partial\mathcal{SMW}_{2}^{2}}{\partial\mathbf{v}_{i}}=2\beta\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{v}_{i}-\mathbf{v}_{j}^{(Q)})=2\beta\biggr{(}p_{i}\mathbf{v}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{v}_{j}^{(Q)}\biggr{)},$
which, for $\beta>0$, is zero if and only if $\mathbf{v}_{i}=T_{\omega}(\mathbf{v}_{i}^{(P)})$. ∎
Remark. In equation 12, we are heuristically adding a label regularization term to the $\mathcal{MW}_{2}$ distance. The actual continuous counterpart (between samples, rather than components) is currently beyond the scope of this paper, but methodologically, this choice remains valid, and is closer to the contributions of [chen2018optimal].